OpenStudy (anonymous):

prove cotx-tanx=(cos2x)/(sinx)(cosx)

OpenStudy (jdoe0001):

is it \[cot(\theta)-tan(\theta)=\frac{cos(2\theta)}{sin(\theta)*cos(\theta)}\]